### v1

# Refined curve counting with descendants and quantum mirrors

Abstract.
We establish a formula for structure constants of the quantum mirror to a log Calabi Yau surface $(Y,D)$ in terms of descendent logarithmic Gromov–Witten invariants of $(Y,D)$. Our result generalises the weak Frobenius structure conjecture for surfaces to the $q$-refined setting, and is proved by relating these invariants to counts of quantum broken lines in the associated quantum scattering diagram.
Introduction
We study the logarithmic Gromov–Witten theory of log Calabi–Yau surfaces. Our results connect enumerative invariants, mirror symmetry, quantization, and quantum scattering diagrams of this surface geometry.
Fix a tuple $m=(m_{1},\dots,m_{n})$ of elements in $\mathbb{Z}^{2}$. Associated to this tuple are two structures: a log Calabi–Yau surface $Y_{m}$, defined in Section 1.1; and a quantum scattering diagram, built from the initial rays $-\mathbb{R}_{\geq 0}m_{i}$ for $i=1,\dots,n$, see Section 2.1. Fix also a tuple $\vartheta=(r_{1},\dots,r_{s})$ of vectors in $\mathbb{Z}^{2}$. Associated to the above data are two sets of enumerative invariants.
  1. (1)
    For any genus $g\geq 0$, and $p_{1},\dots,p_{n}$ integers such that $\sum_{i=1}^{n}p_{i}m_{i}=\sum_{j=1}^{s}r_{j}$, we have the logarithmic Gromov–Witten invariant
    $N_{g,\vartheta}^{\beta_{p}}=\int_{[\mathsf{M}_{g,\vartheta}(Y_{m},\beta_{p})]^{\text{\rm vir}}}(-1)^{g}\lambda_{g}\operatorname{\mathsf{ev}}^{*}(\mathsf{pt})\psi^{s-2}$
    defined as an integral on the moduli space of stable logarithmic maps to $Y_{m}$ of genus $g$ and degree $\beta_{p}$, indicating intersection $p_{i}$ with the exceptional corresponding to $m_{i}$, and tangency given by $\vartheta$.
  2. (2)
    With notation as above we define a combinatorial invariant
    $\langle\hat{\vartheta}_{r_{1}},\dots,\hat{\vartheta}_{r_{s}}\rangle=\sum_{(\gamma_{1},\dots,\gamma_{s})}c(\gamma_{1})\dots c(\gamma_{s})\prod_{2\leq i<j\leq s}q^{\frac{1}{2}s(\gamma_{i})\wedge s(\gamma_{j})}$
    where the sum is over $(\gamma_{1},\dots,\gamma_{s})$ quantum broken lines in the associated quantum scattering diagram with start direction $\vartheta$, meeting at a fixed, balanced point $Q$.
Theorem A.
There is an equality,
$\sum_{p}\left(\sum_{g\geq 0}N_{g,\vartheta}^{\beta_{p}}u^{2g}\right)t_{1}^{p_{1}}\dots t_{n}^{p_{n}}=\frac{1}{(s-1)!}\sum_{\omega\in\Omega_{s}}\langle\hat{\vartheta}_{r_{\omega(1)}},....,\hat{\vartheta}_{r_{\omega(s)}}\rangle.$
Where $p=(p_{1},\dots,p_{n})\in\mathbb{N}^{n}$ such that $\sum_{i=1}^{n}p_{i}m_{i}=\sum_{j=1}^{s}r_{j}$.
Theorem A suggests a method for computing higher genus descendant logarithmic Gromov–Witten invariants of $Y_{m}$. The connection between algebro-geometric invariants and combinatorial invariants is parallel to tropical correspondence theorems for toric targets. Moreover, our results give further evidence for the logarithmic GW/DT correspondence [32] for log Calabi–Yau surfaces, as the form of the answer has rationality properties exactly predicted by the conjecture. If we take the conjecture as true, this gives the first computation of the descendant series of a non-toric target.
Furthermore, our results show that the, a priori mysterious, agreement of descendant counts with structure constants from intrinsic surface mirrors of Gross–Siebert program seem to persist to the $q$-refined setting.
Interpretation in the Gross–Siebert program
The intrinsic mirror construction of the Gross–Siebert program [22, 23] assigns to a log Calabi–Yau pair $(X,D)$ a ring $R$, which is understood to be the ring of functions on the mirror. The authors conjecture  [22, Conjecture 9.2] a connection, the weak Frobenius structure conjecture, between the structure constants of the mirror and the descendant logarithmic Gromov–Witten theory of $(X,D)$. The weak Frobenius structure conjecture has since been proved in complete generality [24], and has had applications [28, 25] to the Fanosearch program [15].
In this surface situation Gross, Hacking, and Keel constructed a mirror equipped with the structure of a formal Poisson variety[19]. The formal Poisson variety structure suggests the existence of a deformation quantization of the mirror algebra, which has subsequently been constructed [11]. The right hand side of Theorem A can be interpreted as structure constants of the quantum mirror.
Let $(Y,D)$ be a Looijenga pair. Associated to this pair is a non-commutative $\mathbb{C}$ algebra given by the quantum mirror to $(Y,D)$. This algebra has a basis given by $\hat{\vartheta}_{r}$ for $r$ an integral point in the tropicalisation $\mathsf{Trop}(Y,D)\cong\mathbb{Z}^{2}$ of $(Y,D)$. Let
$\hat{\vartheta}_{r_{1}},....,\hat{\vartheta}_{r_{s}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}⟨$⟩
denote the coefficient of $1=\hat{\vartheta}_{(0,0)}$ in the algebra product of the basis elements $\hat{\vartheta}_{r_{1}},....,\hat{\vartheta}_{r_{s}}$ [11, Theorem 5.1]. Unlike in the classical setting, this bracket is not symmetric.
Corollary B.
There is an equality, after averaging over cyclic permutations.
$\sum_{\beta}z^{\beta}\left(\sum_{g\geq 0}N_{g,\vartheta}^{\beta}u^{2g}\right)=\frac{1}{(s-1)!}\sum_{\omega\in\Omega_{s}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\hat{\vartheta}_{r_{\omega(1)}},....,\hat{\vartheta}_{r_{\omega(s)}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ .
The proof of Theorem A follows a higher genus version of the strategy of the proof of the weak Frobenius structure conjecture in [29]. We use the technique of factoring the quantum scattering diagram [20], as well as the connection between the Hamiltonians in the consistent quantum scattering diagram and tropical curves [12, 16] to relate the right hand side to tropical curve counts with an $s$-valent vertex. The refined multiplicity of the $s$-valent vertex appears due to the non-commutative product rule of the quantum torus, after symmetrising. We then use the degeneration formula in logarithmic Gromov–Witten theory, together with the refined descendant tropical correspondence theorem of [26] to connect the left hand side to the same tropical curve counts. As in [26], there are extra subtleties in the degeneration formula in this higher genus, descendant setting.
0.1. Future directions
0.1.1. Threefolds
Theorem A can be viewed as a result for the “local log Calabi–Yau surface threefold”, $Y_{m}\times\mathbb{A}^{1}$. The virtual dimension of the moduli space in the case of threefolds is independent of the genus and invariants on $Y_{m}\times\mathbb{A}^{1}$, defined via virtual localisation, are exactly invariants of $Y_{m}$ with a $\lambda_{g}$ insertion. One future direction would be to generalise to the situation of honest threefolds, prompting questions about a higher genus analogue of [5].
0.1.2. K3 surfaces
In a different direction, one could combine Theorem A with the use type III degenerations of K3 surfaces to probe their enumerative geometry. The components of the special fibre of such a degeneration will be of the form $Y_{m}$, for example in the quartic K3 degeneration. The degeneration formula will be more subtle, see  [31, Section 4], due to the presence of the reduced virtual class, particularly in the case of non-primitive classes [8]. However, even in the case of primitive curve classes, if the scattering computations can be understood well, this provides an avenue to produce new calculations. For example, a generalisation of  [31, Theorem 3] with descendants.
0.2. Related work
  • In [13], the authors study five enumerative theories associated to a log Calabi–Yau surface. They propose correspondences relating logarithmic, local and open Gromov–Witten invariants, as well as quiver DT invariants and BPS invariants associated to this log Calabi–Yau surface. In particular, in  [13, Proposition 4.2], a proof is sketched relating higher genus invariants with a $\lambda_{g}$ insertion of log Calabi–Yau surfaces to counts of tuples of three quantum broken lines.
  • In [18], the authors study the geometry of a smooth Fano surface relative to a smooth anticanonical divisor. An interpretation for a certain quantum theta function is given in terms of a $q$-refinement of the open mirror map defined by quantum periods of mirror curves for outer Aganagic–Vafa branes on the associated local threefold. The proof of [18, Proposition 3.12], in particular, relates to Theorem A and  [26, Theorem A].
0.3. Acknowledgements
We thank Sam Johnston, Dhruv Ranganathan, Yannik Schuler and Michel van Garrel for a number of helpful conversations. We thank Tim Gräfnitz for helpful comments on a previous version. P.K.-H. was supported by an Early Career Fellowship at the University of Sheffield awarded by the London Mathematical Society, and Peterhouse, Cambridge. Q.S. was supported by UKRI Future Leaders Fellowship through grant number MR/T01783X/1 and by the starting grant ’Correspondences in enumerative geometry: Hilbert schemes, K3 surfaces and modular forms’, No 101041491 of the European Research Council.
1. Logarithmic Gromov–Witten theory and tropical curves
1.1. Looijenga pairs
The basic geometry in this paper is a Looijenga pair: that is a pair $(Y,D)$ consisting of a smooth complex projective surface $Y$, together with a simple normal crossings, singular, reduced, anticanonical divisor $D$ on $Y$.
Proposition 1.1.1.
[19, Proposition 1.3] Given a Looijenga pair $(Y,D)$ there exists, after potentially corner blow-ups, a toric surface $\overline{Y}_{m}$ and a blow down morphism
$\nu:Y\rightarrow\overline{Y}_{m}$
for $\overline{Y}_{m}$ a toric surface and $\nu$ the blowup in a collection of points in the interior of the toric boundary.
Thus, to study all Looijenga pairs, it suffices to study the geometry described in Setting 1.1.2.
Setting 1.1.2.
Fix $m=(m_{1},\dots,m_{n})$ a tuple of elements of $M=\mathbb{Z}^{2}$. After adding more elements if necessary to $m$ to form a tuple $m^{\prime}$ which spans $\mathbb{R}^{2}$, there is a unique two dimensional fan $\Sigma_{m}$ with rays $-\mathbb{R}_{\geq 0}m_{j}$ for $m_{j}\in m^{\prime}$. We write $\overline{Y}_{m}$ for the toric surface with fan $\Sigma_{m}$. Write $D_{m_{j}}$ for the component of the toric boundary of $Y_{m}$ associated under the toric dictionary to the ray $-\mathbb{R}_{\geq 0}m_{j}$
For each $m_{j}$ for $j=1,\dots,n$, choose a point $x_{j}\in D_{j}$ such that $x_{i}\neq x_{j}$ whenever $i\neq j$ and define a surface $Y_{m}$ to be the blowup $\nu:Y_{m}\rightarrow\overline{Y}_{m}$ at the points $x_{1},\dots,x_{n}$. Let $E_{j}$ denote the exceptional divisor over $x_{j}$ and $\partial Y_{m}$ the strict transform of the toric boundary divisor. Notice $(Y_{m},\partial Y_{m})$ is a Looijenga pair.
In Setting 1.1.2 a choice of tuple $m^{\prime}$ extending $m$ was made. Since logarithmic Gromov–Witten invariants are insensitive to logarithmic modifications  [4, Theorem 1.1.1], our main theorem is insensitive to this choice, and we fix a choice without further comment.
1.2. Logarithmic Gromov–Witten theory
Let $\vartheta=(r_{1},\dots r_{s})$ a tuple of vectors in $M=\mathbb{Z}^{2}$, and replace $m^{\prime}$ by $m^{\prime}~{}\cup~{}\vartheta$, so that every $\mathbb{R}_{\geq 0}r_{i}$ is a ray in the fan of $\overline{Y}_{m}$. We say that a map from an $(s+1)$-pointed curve to ${Y}_{m}$ has tangency given by $\vartheta$ if the $i^{\mathrm{th}}$ marked point hits the strict transform of the toric boundary divisor corresponding to $r_{i}$ with tangency given by the lattice length of $r_{i}$, in other words, the maximal positive integer $\ell$ such that one can write $r_{i}=\ell\overline{r}_{i}$ for $\overline{r}_{i}$ a vector in $M$ called the direction of $r_{i}$.
The moduli space parametrising $(s+1)$-pointed genus $g$ stable maps to $Y_{m}$ with curve class $\beta$ and tangency given by $\vartheta$ at the first $s$ markings, is not proper. The moduli space of stable logarithmic maps $\mathsf{M}_{g,\vartheta}(Y_{m},\beta)$ is a compactification [1, 14, 21].
Writing ¯$\overline{\mathcal{M}}_{g,s+1}$ for the moduli space of stable genus $g$ curves with $s+1$ marked points. This space comes equipped with universal curve $\mathsf{pr}:\mathcal{C}\rightarrow\overline{\mathcal{M}}_{g,s+1}$ and a forgetful morphism
$\pi^{\prime}:\mathsf{M}_{g,\vartheta}(Y_{m},\beta)\rightarrow\overline{\mathcal{M}}_{g,s+1}.$
The moduli space of stable curves carries two flavours of tautological bundle of import to us.
  • The Hodge bundle $\mathbb{E}_{g}=\mathsf{pr}_{\star}\omega_{p}$ where $\omega_{p}$ is the relative dualising sheaf of $p$. We write $\lambda_{g}=c_{g}(\mathbb{E}_{g})$.
  • Note ¯$\overline{\mathcal{M}}_{g,s+1}$ carries $s+1$ tautological sections identifying the marked points. Denote the last section by $S$ and define
    $\psi=c_{1}(S^{\star}\omega_{p}).$
Both classes can be pulled back along $\pi$ to define tautological classes on $\mathsf{M}_{g,\vartheta}(Y_{m},\beta)$ which, by abuse of notation, we also call $\psi$ and $\lambda_{g}$.
We have an evaluation morphism $\operatorname{\mathsf{ev}}:\mathsf{M}_{g,\vartheta}(Y_{m},\beta)\rightarrow Y_{m}$ which evaluates at the $(s+1)^{\mathrm{st}}$ marked point.
1.2.1. Invariants
The moduli space $\mathsf{M}_{g,\vartheta}(Y_{m},\beta)$ carries a virtual fundamental class $\left[\mathsf{M}_{g,\vartheta}(Y_{m},\beta)\right]^{\mathsf{vir}}$ allowing us to define logarithmic Gromov–Witten invariants. We will consider the following descendant logarithmic Gromov–Witten invariants with a $\lambda_{g}$ insertion.
$\mathsf{N}_{g,\vartheta}^{\beta}=\int_{\left[\mathsf{M}_{g,\vartheta}(Y_{m},\beta)\right]^{\mathsf{vir}}}(-1)^{g}\lambda_{g}\operatorname{\mathsf{ev}}^{*}(\mathrm{pt})\psi^{s-2}.$
The classes $\lambda_{g}$ and $\psi$ could instead be defined directly on the moduli space $\mathsf{M}_{g,\vartheta}(Y_{m},\beta)$, call these classes $\overline{\psi}$ and $\overline{\lambda}_{g}$.
Proposition 1.2.1.
$\overline{\lambda}_{g}=\lambda_{g},\quad\quad\mathsf{N}_{g,\vartheta}^{\beta}=\int_{\left[\mathsf{M}_{g,\vartheta}(Y_{m},\beta)\right]^{\mathsf{vir}}}(-1)^{g}\overline{\lambda}_{g}\operatorname{\mathsf{ev}}^{*}(\mathrm{pt})\overline{\psi}^{s-2}.$
Proof.
The first equality follows from  [26, Remark 2.1]. For the second equality, we can use  [17, Example 6.3.4 (a)] to rewrite $\mathsf{N}_{g,\vartheta}^{\beta}$ as an integral over $\mathsf{M}_{g,\vartheta}(Y_{m},\beta,P)$, where we choose a generic point $P$ in the interior where the marking evaluates to. Now the difference $\psi-\overline{\psi}$ on $\mathsf{M}_{g,\vartheta}(Y_{m},\beta,P)$ is supported on the locus of curves where the component containing this marking is destabilised under the forgetful morphism. We claim this locus is empty. By the argument of  [30, Proposition 3.4] it suffices to show that when this component is rational, it contains at least 3 special points. If this component contains another marking then it would also have to contain a node, because there are necessarily other relative markings which need to be distributed on some component. So the only possibility is that this component does not hit the boundary in any point. But then, considering the image of this curve component in $\overline{Y}_{m}$, this is not possible by a dimension count if we generically choose the $x_{i}$. ∎
The class $\beta_{p}$ of a curve is specified by two data
  1. (1)
    The tuple $p=(p_{1},\dots,p_{n})$ recording for each exceptional $E_{i}$ the intersection product $\beta\cdot E_{i}=p_{i}$.
  2. (2)
    The tangency profile $\vartheta$.
In the toric situation we consider more general invariants; consider a balanced multiset $\Delta^{\circ}$ of vectors in $\mathbb{Z}^{2}\setminus\{(0,0)\}$ and non-negative integers $g$ and $n$. As in  [26, Section 2.1] we have associated to this multiset $\mathsf{M}_{g,\Delta}=\mathsf{M}_{g,\Delta}(X_{\Delta},\beta_{\Delta})$, the moduli space of genus $g$, $|\Delta^{\circ}|+n$-pointed stable logarithmic maps to the toric variety $X_{\Delta}$ with tangency given at the last $|\Delta^{\circ}|$ markings by $\Delta^{\circ}$. Fix $\Delta_{F}$ a subset of $\Delta^{\circ}$ and let
$\operatorname{\mathsf{ev}}_{\Delta_{F}}\colon\mathsf{M}_{g,\Delta}\rightarrow(\partial X_{\Delta})^{|\Delta_{F}|}$
denote the evaluation morphism at the toric boundary divisor $\partial X_{\Delta}$ indexed by the elements of $\Delta_{F}$. Let
$\operatorname{\mathsf{ev}}_{i}:\mathsf{M}_{g,\Delta}\rightarrow X_{\Delta}$
denote the evaluation morphism at the non-tangency markings $i=1,\dots,n$. Fix non-negative integers $k_{1},\dots,k_{n}$ such that
$n-1+|\Delta^{\circ}|=2n+\sum_{i=1}^{n}k_{i}+|\Delta_{F}|$
and let
$\mathsf{N}_{g,\Delta,\Delta_{F}}^{\bf{k}}=\int_{[\mathsf{M}_{g,\Delta}]^{\text{\rm vir}}}(-1)^{g}\lambda_{g}\prod_{i=1}^{n}\operatorname{\mathsf{ev}}_{i}^{*}(\mathsf{pt})\psi^{k_{i}}\operatorname{\mathsf{ev}}_{\Delta_{F}}^{*}(r^{|\Delta_{F}|})$
where $r$ is the class of a point on $\partial X_{\Delta}$, and $r^{|\Delta_{F}|}$ denotes the class $\prod_{i}\pi_{i}^{\star}(r)$ in $A^{\star}\left((\partial X_{\Delta})^{|\Delta_{F}|}\right)$.
1.3. Tropical curve counting problems
1.3.1. Tropical curves
We briefly recall notation for tropical curves from [26], see also [10, Section 2.3]. We refer the reader to [2, 30, 33, 34] for background.
Following [3], a graph $\Gamma=(V(\Gamma),E_{f}(\Gamma),E_{\infty}(\Gamma))$ is a triple consisting of finite sets of vertices $V(\Gamma)$, bounded edges $E_{f}(\Gamma)$ and a multiset $E_{\infty}(\Gamma)$ of unbounded edges. We assume all graphs are connected. An abstract tropical curve $|\Gamma|$ is the underlying topological space of a graph $\Gamma$.
Definition 1.3.1.
A parametrised tropical curve ${h}:{\Gamma}\rightarrow\mathbb{R}^{2}$ consists of the following data.
  1. (1)
    A graph $\Gamma$ and a non–negative integer $g_{V}$ assigned to each vertex $V$ of $\Gamma$ called the genus.
  2. (2)
    A bijective function
    $L:E_{\infty}(\Gamma)\rightarrow\{1,...,r+n\}.$
  3. (3)
    A vector weight $v_{V,E}\in\mathbb{Z}^{2}$ for every edge–vertex pair $(V,E)$ with $E\in E_{f}(\Gamma)\cup E_{\infty}(\Gamma)$ and $V\in E$ such that for every vertex $V$, the following balancing condition is satisfied:
    $\sum_{E:V\in E}v_{V,E}=0.$
  4. (4)
    For each bounded edge $E\in E_{f}(\Gamma)$ a positive real number $\ell(E)$, called the length of $E$.
  5. (5)
    A map of topological spaces $h:|\Gamma|\rightarrow\mathbb{R}^{2}$ such that restricting $h$ to the edge $\{v_{1},v_{2}\}$ is affine linear to the line segment connecting $h(V_{1})$ and $h(V_{2})$ and moreover
    $h(V_{2})-h(V_{1})=\ell(E)v_{V_{1},E}.$
    Also, restricting $h$ maps an unbounded edge $E$ associated to vertex $V$ to the ray $h(V)+\mathbb{R}_{\geq 0}v_{V,E}$.
We say $h$ has degree $\Delta$ if $v_{V,E}$ coincides with the $L(E)^{\mathrm{th}}$ column of $\Delta$ whenever $E\in E_{\infty}(\Gamma)$. The genus of a parametrised tropical curve is obtained by adding the sum of $g_{V}$ to the Betti number of $|\Gamma|$. The weight of an edge $E$, denoted $w(E)$, is the lattice length of $v_{V,E}$.
Fixing a subset $\Delta^{F}$ of $\Delta$, and a generic configuration $\{x_{v}\}_{v\in\Delta_{F}}$ of points in $\mathbb{R}^{2}$, we say $h$ has degree $(\Delta,\Delta^{F})$ if it has degree $\Delta$ and the unbounded edges in correspondence with $\Delta^{F}$ asymptotically coincide with the half-lines $x_{v}+\mathbb{R}_{\geq 0}v,v\in\Delta^{F}$.
For a vertex $V$ of $\Gamma$ write $E_{\infty}^{+}(V)$ for the set of unbounded edges $E$ adjacent to $V$ such that $v_{E}\neq 0$ and $E_{f}(V)$ for the set of bounded edges adjacent to $V$. The valency $\mathsf{val}_{V}$ of a vertex $V$ is the cardinality of $E_{f}(V)\cup E_{\infty}^{+}(V)$. Write $\Delta_{V}^{\circ}$ for the multiset of all $v_{V,E}$ for fixed $V$.
1.4. Multiplicities
In the sequel we count parametrised genus zero tropical curves of degree $(\Delta,\Delta^{F})$ satisfying certain incidence conditions. Fix for the remainder of the section a parameterised tropical curve ${h}:\Gamma\rightarrow\mathbb{R}^{2}$. Tropical curves are counted with a multiplicity, closely related to the multiplicity of [6]. This multiplicity is given as a product
$m_{h}(q^{\frac{1}{2}})=\prod_{V\in V(\Gamma)}m_{V}(q^{\frac{1}{2}})$
over multiplicities $m_{V}$ assigned to each vertex $V$ of our tropical curve. We will be counting tropical curves passing through a tuple of points $p=(p_{1},\ldots,p_{n})$ in $\mathbb{R}^{2}$, and thus vertices of $\Gamma$ come in two flavours. A vertex is pointed if its image under $h$ coincides with one of the $p_{i}$. Vertices which are not pointed are unpointed.
We define $v_{1}\wedge v_{2}$ to be the determinant of the matrix with first column $v_{1}$ and second column $v_{2}$. The cyclic group with $m$ elements acts on the set of ordered tuples of $m$ distinct elements from the set $\{1,\ldots,m\}$. The action is induced by sending the integer in position $i$ to position $i+1$ mod $m$. The set of orbits of this action is the set $\Omega_{m}$ of cyclic permutations. For $\omega$ a cyclic permutation, choose an ordered tuple $\tilde{\omega}$ in the orbit $\omega$. Fix a balanced tuple of vectors $(a_{1},\ldots,a_{m})$. Define
$k(\omega)=\sum_{2\leq i<j\leq m}a_{\tilde{\omega}(i)}\wedge a_{\tilde{\omega}(j)}$
where $\tilde{\omega}(i)$ sends $i$ to the element in the $i^{\mathsf{th}}$ position of the chosen representative $\tilde{\omega}$. As the vectors $a_{i}$ have sum zero, $k(\omega)$ is well defined.
Definition 1.4.1.
Let $V$ be a pointed vertex of valency $m$ given by vectors $a_{1},\dots,a_{m}\in\mathbb{Z}^{2}$. The multiplicity of $v$ is
$m_{V}=\frac{1}{(m-1)!}\sum_{\omega\in\Omega_{N}}q^{\frac{k(\omega)}{2}}.$
If however $V$ is an an unpointed trivalent vertex with outgoing vector weights the balanced set of vectors $(a_{1},a_{2},a_{3})$. The multiplicity assigned to $V$ is the Block–Göttsche multiplicity [7].
(1) $m_{V}=\frac{q^{\frac{1}{2}|a_{1}\wedge a_{2}|}-q^{-\frac{1}{2}|a_{1}\wedge a_{2}|}}{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}.$
The multiplicity of a bivalent vertex is $1$.
1.5. Tropical counting problem
We fix once and for all a subset $\Delta^{F}$ of $\Delta$, and a generic configuration $\{x_{v}\}_{v\in\Delta_{F}}$ of points in $\mathbb{R}^{2}$. Thus it will make sense to discuss tropical curves of degree $(\Delta,\Delta^{F})$. For $p$ a generic tuple of points in $\mathbb{R}^{2}$, let $T_{\Delta,\Delta^{F},p}^{\textbf{k}}$ be the set of rigid parametrised tropical curves of degree $(\Delta,\Delta^{\prime})$ passing through $p$ with multiplicity k.
Proposition 1.5.1.
There is an open dense subset $U_{n}^{\textbf{k}}(\Delta_{F})$ of $\mathsf{Ev}^{\mathsf{trop}}\times\mathbb{R}^{2|\Delta_{F}|}$ such that whenever $p,(x_{v})_{v\in\Delta^{F}}\in U_{n}^{\textbf{k}}(\Delta_{F})$ then $T_{\Delta,\Delta^{F},p}^{\textbf{k}}$ is a finite set and the valency of the vertex supporting unbounded edge $E_{i}$ is $k_{i}+2$. Moreover we may choose $U_{n}^{\textbf{k}}(\Delta_{F})$ such that all parametrised tropical curves passing through $p$ with degree k are rigid.
Proof.
The argument of [26, Proposition 1.4.1] works with a cosmetic modification. ∎
In the sequel we assume $p_{i}\neq p_{j}$ whenever $i$ and $j$ are distinct without further comment.
Remark 1.5.2.
Since there are only finitely many combinatorial types of rigid parametrised tropical curves of degree $\Delta$, it is automatic that the set $T_{\Delta,p}^{\textbf{k}}$ is finite.
Recall notation $m_{V}$ for multiplicities of vertex $V$ defined in Section 1.4.
Definition 1.5.3.
Fixing $(p,x)=((p_{1},\ldots,p_{n}),(x_{v})_{v\in\Delta^{F}})$ in $U_{n}^{\textbf{k}}(\Delta_{F})$ and define
$\mathsf{N}^{\Delta,\bf{k}}_{\mathsf{trop},\Delta_{F}}(q)=\sum_{h\in T_{\Delta,\Delta^{\prime},p}^{\textbf{k}}}\prod_{V\in V(\Gamma)}m_{V}(q).$
Define also ${N}_{\mathsf{trop}}^{\Delta,\textbf{k}}(1)={N}_{\mathsf{trop}}^{\Delta,\textbf{k}}$.
A priori the count ${N}_{\mathsf{trop}}^{\Delta,\textbf{k}}(q)$ depends on the choice of point in $U_{n}^{\textbf{k}}(\Delta_{F})$. We suppress this dependence from our notation as it is independent a posteriori.
1.6. The tropical counting and logarithmic Gromov–Witten theory
Fix a subset $\Delta^{F}\subset\Delta$.
Theorem 1.6.1.
After the change of variables $q=e^{iu}$, there is an equality of generating series,
$\sum_{g\geq 0}\mathsf{N}_{g,\Delta,\Delta_{F}}^{\bf{k}}u^{2g-2+|\Delta^{\circ}|-\sum k_{i}}=\prod_{v\in\Delta^{F}}\frac{1}{|v|}\mathsf{N}^{\Delta,\bf{k}}_{\mathsf{trop},\Delta_{F}}(q)\left((-i)(q^{\frac{1}{2}}-q^{-\frac{1}{2}})\right)^{|\Delta^{\circ}|-\sum_{i}k_{i}-2}.$
Proof.
The theorem is almost identical to  [26, Theorem A]. Indeed, applying the degeneration formula yields the analogue of  [26, Proposition 3.4.2]. In absorbing the contribution of weights of bounded edges as in  [26, Section 6.1], one must divide through by the weights of the fixed unbounded edges to compensate, which explains the prefactor $|v|^{-1}$ of the right-hand side the statement. Since the vertex contributions are unaffected the theorem then follows from the vertex contribution calculations in  [26, Section 6.2]. ∎
2. Scattering diagrams and tropical curves
Let $M=\operatorname{Hom}(T,\mathbb{C}^{*})\cong\mathbb{Z}^{2}$ be the character lattice of a two-dimensional torus $T$. We will make use of the ‘algebra of functions’ of the quantum torus
$\mathbb{C}_{q}[M]=\bigoplus_{m\in M}\mathbb{C}[q^{\pm\frac{1}{2}}]\hat{z}^{m},$
the $\mathbb{C}[q^{\pm\frac{1}{2}}]$ algebra with non-commutative product given by $\hat{z}^{r}\cdot\hat{z}^{r^{\prime}}=q^{\frac{1}{2}r\wedge r^{\prime}}\hat{z}^{r+r^{\prime}}$ for $r,r^{\prime}\in M$, as well as the change of variables $q=e^{iu}$.
2.1. Quantum scattering diagrams
Let $R$ be a complete local $\mathbb{C}$ algebra with maximal ideal $\mathbf{m}_{R}$. A quantum scattering diagram over a ring $R$ is the data of a collection of half lines $\{\mathfrak{d}\}$ called rays in $M_{\mathbb{R}}\cong\mathbb{R}^{2}$ with primitive integral direction $m_{\mathfrak{d}}\in M\setminus\{0\}$. To each ray is an associated function called the Hamiltonian $H_{\mathfrak{d}}$ such that
[…]
See [12, Section 1] for further details on quantum scattering diagrams.
The Looijenga pair $(Y_{m},\partial Y_{m})$ has an associated quantum scattering diagram called the initial quantum scattering diagram.
Definition 2.1.1.
The initial quantum scattering diagram $\hat{\mathfrak{D}}_{\mathfrak{m}}$ associated to the tuple $m$ is the quantum scattering diagram over $R:=\mathbb{C}\llbracket t_{1},...,t_{n}\rrbracket$ consisting of incoming rays $(\mathfrak{d}_{j},\hat{H}_{\mathfrak{d}_{j}})$ for $1\leqslant j\leqslant n$, where
$\mathfrak{d}_{j}=-\mathbb{R}_{\geqslant 0}m_{j}\,,$
and
(2) $\hat{H}_{\mathfrak{d}_{j}}=\sum_{\ell\geqslant 1}\frac{1}{\ell}\frac{(-1)^{\ell-1}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}t_{j}^{\ell}\hat{z}^{\ell m_{j}}\,,$
where $q=e^{iu}$.
Associated to any scattering diagram is its consistent completion, see  [27, Theorem 6].
Definition 2.1.2.
Let $S(\hat{\mathfrak{D}}_{\mathfrak{m}})$ be the quantum scattering diagram obtained by forming the consistent completion of $\hat{\mathfrak{D}}_{\mathfrak{m}}$, as in [27].
Theorem A relates logarithmic Gromov–Witten invariants of $(Y_{m},\partial Y_{m})$ to quantum broken line counts on $S(\hat{\mathfrak{D}}_{\mathfrak{m}})$, see Section 2.4. These counts also determine the product structure on the quantum mirror algebra to $(Y_{m},\partial Y_{m})$, see Section 2.5.
2.2. Factoring the quantum scattering diagram
The consistent quantum scattering diagram $S(\hat{\mathfrak{D}}_{\mathfrak{m}})$ is combinatorially complex, but can be understood in terms of factored quantum scattering diagrams whose structure reflects the combinatorics of tropical curves. This connection was established in [16]. We follow  [12, Section 6].
To factor a scattering diagram over the ring $R=\mathbb{C}\llbracket t_{1},...,t_{n}\rrbracket$, we study a scattering diagram over the ring $R_{N}=\mathbb{C}[t_{1},\dots,t_{n}]/(t_{1}^{N+1},\dots,t_{n}^{N+1})$. There is an embedding
$R_{N}\hookrightarrow\tilde{R}_{N}=\mathbb{C}[u_{ij}:1\leq i\leq n,1\leq j\leq N]/\langle u_{ij}^{2}:1\leq i\leq n,1\leq j\leq N\rangle$
by
$t_{i}=\sum_{j=1}^{N}u_{ij},\,\,\,\Rightarrow\,\,\,t^{\ell}_{i}=\sum_{A\subset\{1,\dots,N\},|A|=\ell}\ell!\prod_{a\in A}u_{ia}.$
Under this embedding we have
$\hat{H}_{\mathfrak{d}_{j}}=\sum_{\ell=1}^{N}\sum_{A\subset\{1,\dots,N\},|A|=\ell}\left(\frac{1}{\ell}\frac{(-1)^{\ell-1}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)\ell!\left(\prod_{a\in A}u_{ja}\right)\hat{z}^{\ell m_{j}}$
We then define a new initial quantum scattering diagram $\hat{\mathfrak{D}}_{\mathfrak{m}}^{\mathrm{split}}$ with initial rays $(\mathfrak{d}_{j\ell A},\hat{H}_{j\ell A})$ for $\ell=1,\dots,N,A\subset\{1,\dots,N\},|A|=\ell$, where
$\mathfrak{d}_{j\ell A}=-\mathbb{R}_{\geq 0}m_{j}+c_{j\ell A}$
for some generic $c_{j\ell A}\in\mathbb{R}^{2}$ in general position, and
$\hat{H}_{\mathfrak{d}_{j\ell A}}=\left(\frac{1}{\ell}\frac{(-1)^{\ell-1}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)\ell!\left(\prod_{a\in A}u_{ja}\right)\hat{z}^{\ell m_{j}}.$
By  [12, Lemma 6.1], the Hamiltonian attached to any ray in $S(\hat{\mathfrak{D}}_{\mathfrak{m}}^{\mathrm{split}})$ can be determined recursively from the the initial rays and Hamiltonians.
2.3. From factored scattering diagrams to tropical curves
Lemma 2.3.1.
[12, Proof of Proposition 6.2] For a general choice of $\hat{\mathfrak{D}}_{\mathfrak{m}}^{\mathrm{split}}$ there is a bijective correspondence between $(\mathfrak{d},H_{\mathfrak{d}})\in S(\hat{\mathfrak{D}}_{\mathfrak{m}}^{\mathrm{split}})$, and rational tropical curves $h:\Gamma\rightarrow\mathbb{R}^{2}$ such that
  1. (i)
    There is an edge $E_{\mathsf{out}}\in E_{\infty}(\Gamma)$ with $h(E_{\mathsf{out}})=\mathfrak{d}$.
  2. (ii)
    If $E\in E_{\infty}(\Gamma)\setminus\{E_{\mathsf{out}}\}$ or if $E_{\mathsf{out}}$ is the only edge of $\Gamma$ and $E=E_{\mathsf{out}}$, then $h(E)$ is contained in some $\mathfrak{d}_{j\ell A}$ where
    $1\leq j\leq n,\ \ \ A\subseteq\{1,\ldots,N\},\ \ \ \ell\geq 1$
    Furthermore, if $E\not=E_{\mathsf{out}}$, the unbounded direction of $h(E)$ is given by $-m_{i}$.
  3. (iii)
    If $E,E^{\prime}\in E_{\infty}(\Gamma)\setminus\{E_{\mathsf{out}}\}$ and $h(E)\subseteq\mathfrak{d}_{j\ell A}$ and $h(E^{\prime})\subseteq\mathfrak{d}_{j\ell^{\prime}A^{\prime}}$, then $A\cap A^{\prime}=\emptyset$.
  4. (iv)
    If $E\in E_{\infty}(\Gamma)\setminus\{E_{\mathsf{out}}\}$ or if $E_{\mathsf{out}}$ is the only edge of $\Gamma$ and $E=E_{\mathsf{out}}$, and $h(E)\subseteq\mathfrak{d}_{j\ell A}$, we have $w_{\Gamma}(E)=\ell$.
Furthermore we have that
(3) $\hat{H}_{\mathfrak{d}}=m_{\Gamma}(q^{\frac{1}{2}})\prod_{j=1}^{n}\prod_{\ell\geq 1}\left(\frac{(-1)^{\ell}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell,j}}(\ell!)^{k_{\ell,j}}\left(\prod_{A\in A^{\Gamma}_{j\ell}}\prod_{a\in A}u_{ja}\right)\frac{\hat{z}^{\ell_{p}m}}{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}$
where $m_{\Gamma}(q^{\frac{1}{2}})$ is the refined multiplicity of the tropical curve $h:\Gamma\rightarrow\mathbb{R}^{2}$, $A_{j\ell}^{\Gamma}$ is the set of subsets of $\{1,\dots,N\}$ of size $\ell$ such that $\mathfrak{d}_{j\ell A}$ is an unbounded ray of $h:\Gamma\rightarrow\mathbb{R}^{2}$, and $k_{\ell j}=|A_{j\ell}^{\Gamma}|$.
2.4. Monomials on broken lines
Definition 2.4.1.
A quantum broken line is a pair $(\gamma,\{m_{L}\})$ as follows.
  1. (1)
    A proper continous piecewise integral affine map
    $\gamma:(-\infty,0]\rightarrow M_{\mathbb{R}}$
    with only finitely many domains of linearity.
  2. (2)
    For each $L\subset(-\infty,0]$ a maximal connected domain of linearity of $\gamma$, a choice of monomial $m_{L}=c_{L}\hat{z}^{p_{L}}$ with $c_{L}\in R_{u}=R\llbracket u\rrbracket$, $p_{L}\in M$.
Where these data satisfy the following
  • along any domain of linearity the direction is given by $-p_{L}$
  • $\gamma(0)=Q\in M_{\mathbb{R}}$
  • For the unique unbounded domain of linearity $L$, $m_{L}=\hat{z}^{r}$ for some $r\in M\setminus\{0\}$.
  • Let $t\in(-\infty,0)$ be a point of $\gamma$ which is not linear, passing from the domain of linearity $L$ to the domain of linearity $L^{\prime}$. Then $\gamma$ necessarily crosses a ray $\mathfrak{d}$ in the scattering diagram at time $t$. Moreover $c_{L^{\prime}}\hat{z}^{p_{L^{\prime}}}$ is a summand of the image of $c_{L}\hat{z}^{p_{L}}$ under the automorphism
    $c_{L}\hat{z}^{p_{L}}\mapsto\exp(\hat{H}_{\mathfrak{d}})c_{L}\hat{z}^{p_{L}}\exp(-\hat{H}_{\mathfrak{d}})$
We denote the monomial attached to the segment ending at $Q$ by $c(\gamma)\hat{z}^{s(\gamma)}$.
Definition 2.4.2.
Let $\vartheta=(r_{1},\dots,r_{s})$ be a tuple of vectors in $M=\mathbb{Z}^{2}$. Define
(4) $\langle\hat{\vartheta}_{r_{1}},\dots,\hat{\vartheta}_{r_{s}}\rangle=\sum_{(\gamma_{1},\dots,\gamma_{s})}c(\gamma_{1})\dots c(\gamma_{s})\prod_{2\leq i<j\leq s}q^{\frac{1}{2}s(\gamma_{i})\wedge s(\gamma_{j})}$
where the sum is over $(\gamma_{1},\dots,\gamma_{s})$ tuples of broken lines such that $\gamma_{i}(0)=Q$, for some fixed $Q$, $m_{L}=\hat{z}^{r_{i}}$, and $\sum_{i=1}^{s}s(\gamma_{i})=0$.
The notation $\hat{\vartheta}_{r_{i}}$ comes from the connection to the product structure on the quantum mirror. Let $R=\mathbb{C}\llbracket t_{1},\dots,t_{n}\rrbracket$.
2.5. Product structure
Let
$R_{(Y_{m},\partial Y_{m})}=\bigoplus_{r\in M}R_{u}\hat{\vartheta}_{r}$
be the free $R_{u}$-module. We call the basis elements $\hat{\vartheta}_{r}$ quantum theta functions. We explain briefly how (4) relates to a product structure on $R_{(Y_{m},\partial Y_{m})}$. Fix a point $Q\in\mathbb{R}^{2}\setminus S(\hat{\mathfrak{D}}_{\mathfrak{m}})$. Let
$\hat{\vartheta}_{r}^{Q}=\sum_{\gamma}c(\gamma)\hat{z}^{s(\gamma)}\in R_{u}\otimes_{\mathbb{C}[q^{\pm\frac{1}{2}}]}\mathbb{C}_{q}[M]$
where the sum is over all broken lines $\gamma$ with start direction given by $r$, which end at the point $Q$. We define a product structure on $R_{(Y_{m},\partial Y_{m})}$ via the structure constants
$\hat{\vartheta}_{r_{1}}\dots\hat{\vartheta}_{r_{s}}=\sum_{r}c^{r}_{r_{1}\dots r_{s}}\hat{\vartheta}_{r}$
where $c^{r}_{r_{1}\dots r_{s}}$ is given by the coefficient of $\hat{\vartheta}_{r}^{Q}$ in $\hat{\vartheta}^{Q}_{r_{1}}\dots\hat{\vartheta}^{Q}_{r_{s}}$ for $Q$ sufficiently close to $r$.
Lemma 2.5.1.
$c^{0}_{r_{1}\dots r_{s}}=\langle\hat{\vartheta}_{r_{1}},\dots,\hat{\vartheta}_{r_{s}}\rangle$
Proof.
Taking the product of theta functions gives
$\sum_{(\gamma_{1},\dots,\gamma_{s})}c(\gamma_{1})\hat{z}^{s(\gamma_{1})}\dots c(\gamma_{s})\hat{z}^{s(\gamma_{s})}$
where the sum is over $(\gamma_{1},\dots,\gamma_{s})$ tuples of broken lines such that $\gamma_{i}(0)=Q$, for some fixed $Q$ (near the origin). Taking the coefficient of $\hat{\vartheta}_{0}=1$ ensures that $s(\gamma_{1})+\dots+s(\gamma_{s})=0$ so that we have
$c^{0}_{r_{1}\dots r_{s}}=\sum_{(\gamma_{1},\dots,\gamma_{s})}c(\gamma_{1})\dots c(\gamma_{s})\hat{z}^{-\sum_{i=2}^{s}s(\gamma_{i})}\hat{z}^{s(\gamma_{2})}\dots\hat{z}^{s(\gamma_{s})}$
and the result follows now from product rule $\hat{z}^{r}\cdot\hat{z}^{r^{\prime}}=q^{\frac{1}{2}r\wedge r^{\prime}}\hat{z}^{r+r^{\prime}}$. ∎
Remark 2.5.2.
The quantum mirror of [11] is defined with respect to a different quantum scattering diagram, analogous to the canonical scattering diagram of [19], but the output is essentially the same, and we explain this briefly. The algebra is defined by fixing the data $(B,\Sigma),P,J$ and $\varphi$. Here $(B,\Sigma)$ is an integral affine manifold (with singularities) defined via the geometry of $(Y_{m},\partial Y_{m})$, $P$ is a toric monoid containing $\mathrm{NE}(Y_{m})$, $\varphi$ is a $P_{\mathbb{R}}^{\mathrm{gp}}$-valued PL function on $B$ and $J$ is a monomial ideal of $P$. The product structure is defined also defined with respect to the scattering diagram by counts of broken lines. By [11, Proposition 4.10], the canonical scattering diagram is related to the scattering diagram of Definition 2.1.2, by a piecewise linear map. As a result broken lines go to broken lines and the difference in their definition of the product of [11, Theorem 3.11], cancels with the presence of the PL function to recover the same product structure.
2.6. Quantum scattering to $q$-refined tropical curves
Let
$[\langle\hat{\vartheta}_{r_{1}},\dots,\hat{\vartheta}_{r_{n}}\rangle]^{\mathsf{sym}}=\frac{1}{(s-1)!}\sum_{\omega\in\Omega_{s}}\langle\hat{\vartheta}_{r_{\omega(1)}},....,\hat{\vartheta}_{r_{\omega(s)}}\rangle$
Lemma 2.6.1.
In the factored scattering diagram there is an equality
$[\langle\hat{\vartheta}_{r_{1}},\dots,\hat{\vartheta}_{r_{n}}\rangle]^{\mathsf{sym}}=\sum_{p}\sum_{k\vdash p}\sum_{A_{j\ell}}\left(\sum_{\Gamma\in T_{A_{j\ell}}}\prod_{V\in V(\Gamma)}m_{V}(q^{\frac{1}{2}})\right)\prod_{j=1}^{n}\prod_{\ell\geq 1}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell j}}\left(\ell!^{k_{\ell j}}\right)\left(\prod_{A\in A_{j\ell}}\prod_{a\in A}u_{ja}\right)$
The first sum is over all $p=(p_{1},\dots,p_{n})$ such that $\sum_{i=1}^{n}p_{i}m_{i}=\sum_{j=1}^{s}r_{j}$. Here, $k\vdash p$ denotes a partition for each $j=1,\dots,n$ $(k_{\ell j})_{\ell\geq 1}$ of $p_{j}$, $A_{j\ell}$ is a set of $k_{\ell j}$ disjoint subsets of $\{1,\dots,N\}$ of size $\ell$ and $T_{A_{j\ell}}$ is the set of genus-zero tropical curves $\Gamma$ having an unbounded edge of asymptotic direction $r_{i}$ for $i=1,\dots,s$ as well as for every $j=1,\dots,n,\ell\geq 1,A\in A_{j\ell}$ an unbounded edge of weight $\ell$ coniciding with $\mathfrak{d}_{j\ell A}$, with an $s$-valent point condition at a point $Q\in\mathbb{R}^{2}$.
Proof.
We have that
(5) $\langle\vartheta_{r_{1}},\dots,\vartheta_{r_{n}}\rangle=\sum_{(\gamma_{1},\dots,\gamma_{s})}c(\gamma_{1})\cdot\dots c(\gamma_{s})\prod_{2\leq i<j\leq s}q^{\frac{1}{2}s(\gamma_{i})\wedge s(\gamma_{j})}$
where the sum is taken over tuples of quantum broken lines where $\gamma_{i}$ has start direction $r_{i}$, ends at $Q$ and such that the broken lines collectively are balanced at $Q$. For each $i=1,\dots,s$, $\gamma_{i}$ bends at various rays $\mathfrak{d}$ in the factored scattering diagram. Recall by Lemma 2.3.1 there is a bijection between rays $\mathfrak{d}$ of the factored scattering diagram and tropical curves such that the Hamiltonian associated to the ray is given by (3).
If $\hat{z}^{k}$ is the monomial attached to a segment of the broken line $\gamma_{i}$, the wall-crossing automorphism of $\mathfrak{d}$ acts by sending
$\hat{z}^{k}\mapsto\exp(\hat{H_{\mathfrak{d}}})\hat{z}^{k}\exp(-\hat{H}_{\mathfrak{d}})$
Since the variables $u_{ij}^{2}=0$ this implies that if $\gamma_{i}$ bends when it hits $\mathfrak{d}$ at $V$, the outgoing Hamiltonian picks up a factor of
$m_{V}(q^{\frac{1}{2}})\cdot m_{\Gamma}(q^{\frac{1}{2}})\prod_{j=1}^{n}\prod_{\ell\geq 1}\left(\frac{(-1)^{\ell}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell,j}}(\ell!)^{k_{\ell,j}}\left(\prod_{A\in A^{\Gamma}_{j\ell}}\prod_{a\in A}u_{ja}\right)$
where the edges are considered to be the incoming and outgoing edges of the broken line together with the part of $\mathfrak{d}$ which has finite length. Adjoining this part of the tropical curve to $\gamma_{i}$ one can associate to the tuple $(\gamma_{1},\dots,\gamma_{n})$ a tropical curve with an $s$-valent vertex.
On the other hand, every tropical curve with an $s$-valent vertex at $Q$, with asymptotic directions given by $r_{1},\dots,r_{s}$ as well as some subset of the $\mathfrak{d}_{j\ell A}$, gives rise to broken lines $(\gamma_{1},\dots,\gamma_{s})$ appearing in the right hand side of (5). The result now follows from the fact that the $q$-refined multiplicity of the $s$-valent vertex as in Definition 1.4.1, appears from the factor
$\prod_{2\leq i<j\leq s}q^{\frac{1}{2}s(\gamma_{i})\wedge s(\gamma_{j})}$
after averaging over cyclic permutations. ∎
Corollary 2.6.2.
There is an equality
$[\langle\hat{\vartheta}_{r_{1}},....,\hat{\vartheta}_{r_{s}}\rangle]^{\mathsf{sym}}=\sum_{p}\sum_{k\vdash p}\mathsf{N}^{\Delta,s}_{\mathsf{trop,\Delta_{k}}}(q)\left(\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{1}{k_{\ell j}!}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell j}}\right)\left(\prod_{j=1}^{n}t_{j}^{p_{j}}\right)$
The first sum is over all $p=(p_{1},\dots,p_{n})$ such that $\sum_{i=1}^{n}p_{i}m_{i}=\sum_{j=1}^{s}r_{j}$. Here, $k\vdash p$ denotes a partition for each $j=1,\dots,n$ $(k_{\ell j})_{\ell\geq 1}$ of $p_{j}$, $\Delta$ consists of $k_{\ell j}$ copies of the vector $-\ell m_{j}$ as well as $r_{1},\dots,r_{s}$ and $\Delta_{k}$ denotes $\Delta\setminus\{r_{1},\dots,r_{s}\}$.
Proof.
This follows the same argument as in  [12, Proof of Proposition 6.2]. Let $B=\cup_{A\in A_{j\ell}}A$, a subset of $\{1,\dots,N\}$ of size $\sum_{\ell\geq 1}\ell k_{\ell j}=p_{j}$. Conversely, the number of ways to write a set $B$ of $p_{j}=\sum_{\ell\geq 1}\ell k_{\ell j}$ elements as a disjoint union of subsets, $k_{\ell j}$ of them being of size $\ell$, is equal to
$\frac{p_{j}!}{\prod_{\ell\geq 1}k_{\ell j}!(\ell!)^{k_{\ell j}}}.$
Replacing the sum over $A_{j\ell}$ by a sum over $B$ and using the fact that
$t_{j}^{p_{j}}=\sum_{B\subset\{1,\dots,N\},|B|=p_{j}}p_{j}!\prod_{b\in B}u_{jb}$
gives the result after identifying
$\mathsf{N}^{\Delta,s}_{\mathsf{trop,\Delta_{k}}}(q)=\sum_{\Gamma\in T_{A_{j\ell}}}\prod_{V\in V(\Gamma)}m_{V}(q^{\frac{1}{2}}).$
3. From logarithmic Gromov–Witten invariants to tropical curve counts
In this section we use the degeneration formula to relate the logarithmic Gromov–Witten invariants of $Y_{m}$ to toric logarithmic Gromov–Witten invariants. The latter can be connected with refined tropical curve counting due to Theorem 1.6.1. This section follows  [11, Section 5].
3.1. Degeneration
First we recall the degeneration of  [20, Section 5.3] (see also [12, 29]). Recall that the tuple $m=(m_{1}\dots,m_{n})$ of non-zero vectors in $\mathbb{Z}^{2}$ fixes a toric surface $\overline{Y}_{m}$. For every $j=1,\dots,n$ let $x_{j}$ denote a general point on the interior of the boundary divisor $D_{m_{j}}$. Take the total space of the deformation to the normal cone of $D_{m_{1}}\cup\dots\cup D_{m_{n}}$ in $\overline{Y}_{m}$ (we can assume the $D_{m_{j}}$ are disjoint). Then blow-up the strict transforms of the constant sections determined by $x_{j}$ to obtain a family
$\epsilon:\mathcal{Y}_{m}\rightarrow\mathbb{A}^{1}.$
The general fibre is $Y_{m}$ and the central fibre is
$\overline{Y}_{m}\cup\bigcup_{j=1}^{n}\tilde{\mathbb{P}}_{j}$
where ~$\tilde{\mathbb{P}}_{j}$ is the blow-up of the projective completion of the normal bundle of $D_{j}$ in $\overline{Y}_{m}$ at the points corresponding to the $x_{j}$. Equipping $\mathcal{Y}_{m}$ with the divisorial logarithmic structure with respect to the central fibre together with the strict transform of $\partial\overline{Y}_{m}\times\mathbb{A}^{1}$ and equipping $\mathbb{A}^{1}$ with the divisorial logarithmic structure with respect to the origin makes $\epsilon$ into a logarithmically smooth morphism. Restricting to the central fibre there is a logarithmically smooth morphism $\mathcal{Y}_{m,0}\rightarrow\mathsf{pt}_{\mathbb{N}}$. Write $\mathsf{M}_{g,\vartheta}(\mathcal{Y}_{m,0},\beta_{p})$ for the moduli space of genus $g$, $(s+1)$-pointed stable logarithmic maps to $\mathcal{Y}_{m,0}$ over $\mathrm{pt}_{\mathbb{N}}$, of class $\beta_{p}$, with tangency given by $\vartheta$.
By deformation invariance of logarithmic Gromov–Witten invariants we have that
$\mathsf{N}_{g,\vartheta}^{\beta_{p}}=\int_{\left[\mathsf{M}_{g,\vartheta}(\mathcal{Y}_{m,0},\beta_{p})\right]^{\mathsf{vir}}}(-1)^{g}\lambda_{g}\operatorname{\mathsf{ev}}^{*}(\mathrm{pt})\psi^{s-2}.$
3.2. Decomposition
Tropical data associated to the morphism $\mathcal{Y}_{m,0}\rightarrow\mathsf{pt}_{\mathbb{N}}$ may be recorded as a polyhedral complex $\mathcal{Y}^{\mathsf{trop}}_{m,0}$ with edge lengths metrised [12, Section 5.3]. Our polyhedral complex is obtained by subdividing the fan of $\overline{Y}_{m}$, whose vertex is called $v_{0}$, by placing a vertex $v_{j}$ at distance 1 along each ray in direciton $m_{j}$ and adding unbounded edges attached to $v_{j}$ parallel to those rays adjacent to $-\mathbb{R}_{\geq 0}m_{j}$.
Figure 1. A example of the central fibre $\mathcal{Y}_{m,0}$ and the tropicalisation $\mathcal{Y}_{m,0}^{\mathsf{trop}}$
The decomposition formula of [2] allows us to write $N_{g,\vartheta}^{\beta_{p}}$ as a sum over rigid tropical curves $h:\Gamma\rightarrow\mathcal{Y}^{\mathsf{trop}}_{m,0}$.
The moduli space of $(s+1)$-pointed genus $g$ stable logarithmic maps marked by $h$, denoted $\mathsf{M}^{h}_{g,\vartheta}$ is a proper Deligne–Mumford stack with a natural perfect obstruction theory and a forgetful morphism
$i_{h}:\mathsf{M}^{h}_{g,\vartheta}\rightarrow\mathsf{M}_{g,\vartheta}(Y_{m,0},\beta_{p}).$
The decomposition formula [2] tells us that
Proposition 3.2.1.
$\mathsf{N}^{\beta_{p}}_{g,\vartheta}=\sum_{h}\frac{n_{h}}{|\mathsf{Aut}(h)|}\mathsf{N}_{h}$
where
$\mathsf{N}_{h}=\int_{\left[\mathsf{M}_{g,\vartheta}^{h}\right]^{\mathsf{vir}}}(-1)^{g}\lambda_{g}\operatorname{\mathsf{ev}}^{*}(\mathrm{pt})\psi^{s-2}.$
In this situation only certain rigid tropical types will have a non-vanishing contribution in the decomposition formula. In order to index these we let $k=(k_{1},\dots,k_{n})$ be a partition of $p=(p_{1},\dots,p_{n})$ i.e. $k_{j}=(k_{\ell j})_{\ell\geq 1}$, finitely many non-zero integers with
$\sum_{\ell\geq 1}\ell k_{\ell j}=p_{j}.$
We write $k\vdash p$. Moreover, let $s(k)=\sum_{j=1}^{n}\sum_{\ell\geq 1}k_{\ell j}$ and $w(k)=(w_{1}(k),\dots,w_{s(k)}(k))$ where each $w_{j}(k)\in\mathbb{Z}^{2}$ and so that $w(k)$ contains $k_{\ell j}$ copies of the vector $-\ell m_{j}$.
Recall that the vectors, $\vartheta_{1},\dots,\vartheta_{s}$, are multiples of the rays of the fan of $\overline{Y}_{m}$. Suppose $\vartheta_{i_{1}},\dots,\vartheta_{i_{b}}$ are multiples of $-m_{i_{1}},\dots,-m_{i_{b}}$.
Definition 3.2.2.
We now define a class of rigid tropical curves
$h_{k,\underline{g}}^{s}:\Gamma_{k,\underline{g}}^{s}\rightarrow\mathcal{Y}_{m,0}^{\mathsf{trop}}$
depending on a partition $k\vdash p$ as well as a $\underline{g}=(g_{0},\dots,g_{s(k)})$ a tuple of non-negative integers, with $\sum_{i=0}^{s(k)}g_{i}=g$.
Definition 3.2.3.
Let $\Gamma_{k,\underline{g}}^{s}$ be the genus zero graph with
  • $V(\Gamma_{k,\underline{g}}^{s})=\{V_{0},\dots,V_{s(k)},V^{\prime}_{1},\dots,V^{\prime}_{b}\}$.
  • $E_{f}(\Gamma_{k,\underline{g}}^{s})=\{E_{1},\dots,E_{s(k)},E^{\prime}_{1},\dots,E^{\prime}_{b}\}$ with $V_{0},V_{j}\in E_{j}$ and $V_{0},V^{\prime}_{j}\in E^{\prime}_{j}$.
  • $E_{\infty}(\Gamma_{k,\underline{g}}^{s})=\{F_{1},\dots,F_{s},F$}, with $V^{\prime}_{j}\in F_{i_{j}}$ for $j=1,\dots,b$, $V_{j}\in F_{j}$ for other $j$, and $V_{0}\in F$.
To the graph $\Gamma^{s}_{k,\underline{g}}$ assign
  • the genus $g_{j}$ to $V_{j}$ for $j=0,\dots,s(k)$ and genus zero to $V^{\prime}_{j}$, $j=1,\dots,b$.
  • the length $\ell(E_{j})=\frac{1}{|w_{j}(k)|}$ for $j=1\dots,s(k)$ and the length $\ell(E^{\prime}_{j})=\frac{1}{|\vartheta_{i_{j}}|}$ for $j=1,\dots,b$.
We define a tropical curve $h_{k,\underline{g}}^{s}:\Gamma^{s}_{k,\underline{g}}\rightarrow\mathcal{Y}^{\mathsf{trop}}_{m,0}$ by
  • $h(V_{0})=v_{0}$ and $h(V_{j})=v_{i}$ if $w_{j}=-\ell m_{i}$, for $j=1,\dots,s(k)$
  • $h(V^{\prime}_{j})=v_{i_{j}}$ and $h(F_{j})=v_{i_{j}}+\mathbb{R}_{\geq 0}\vartheta_{j}$
  • $h(F)=v_{0}$
  • decorate $V_{0}$ with the curve class $\beta_{w(k)}\in H_{2}(\overline{Y}_{m},\mathbb{Z})$, determined by the contact data $w(k)$
  • if $w_{j}=-\ell m_{i}$, decorate $V_{j}$ with the curve class $\ell[C_{i}]$ where $[C_{i}]$ corresponds to the strict transform of a fibre class passing through the point corresponding to $x_{i}$ in ~$\tilde{\mathbb{P}}_{i}$, for $j=1,\dots,s(k)$
  • decorate $V^{\prime}_{j}$ with the curve class $|\vartheta_{i_{j}}|[C^{\prime}_{i_{j}}]$ where $[C^{\prime}_{i_{j}}]$ corresponds to the strict transform of a fibre class in ~$\tilde{\mathbb{P}}_{i_{j}}$, for $j=1,\dots,b$.
Figure 2. Example of $\Gamma_{k,\underline{g}}^{s}$ for target as in Figure 1.
We claim the only rigid tropical types appearing with non-vanishing contribution in the decomposition formula are $h_{k,\underline{g}}^{s}$.
Proposition 3.2.4.
Every rigid tropical type $h$ with a non-zero contribution to the decomposition formula is of the form $h=h^{s}_{k,\underline{g}}$.
The proof will use the following lemmas. We will say that an edge $E$ of a tropical curve is on a 00 line if the line containing $E$ passes through the origin.
Lemma 3.2.5.
Suppose $h\colon\Gamma\rightarrow\mathcal{Y}_{m,0}^{\mathsf{trop}}$ is a rigid tropical type. Fix an edge $\gamma_{0}$, and a vertex $V$ of $\gamma_{0}$. Assume $\gamma_{0}$ is not on a 00 line. Then there is a sequence of edges $\gamma_{0},...,\gamma_{n}$ satisfying the following properties,
  1. (1)
    $\gamma_{i},\gamma_{i+1}$ meet at a vertex, and $\gamma_{0},\gamma_{1}$ meet at $V$.
  2. (2)
    Either $\gamma_{n}$ is unbounded or $\gamma_{n}=\gamma_{0}$
Proof.
We first show that fixing an edge $\gamma_{0}$ not on a zero line, and a vertex $v$ of $\gamma_{0}$, there exists an edge $\gamma_{1}$ which meets $\gamma_{0}$ at $v$ and is not on a zero line. Indeed, this is true by applying the balancing condition at $v$ in the direction e orthogonal to the line joining $v$ to the origin. Since $\gamma_{0}$ is not orthogonal to e, $v$ must be adjacent to some other edge $\gamma_{1}$ whose dot product with e is non-zero.
We now recursively construct our sequence of edges. By the previous paragraph, given a sequence $\gamma_{1},...,\gamma_{i}$ of edges satisfying (1), and such that $\gamma_{i}$ is bounded, there exists an edge $\gamma_{i+1}$ which does not lie on a zero line and which meets $\gamma_{i}$ at the vertex not adjacent to $\gamma_{i-1}$. Tropical curves have finitely many edges, and the process must terminate, which can only occur when we find some unbounded $\gamma_{n}$ or else we have constructed a cycle. ∎
We define a continuable edge to be an edge which is either an unbounded edge, or a bounded edge connected to a 2-valent vertex whose other edge is unbounded.
Lemma 3.2.6.
Suppose a rigid type $h\colon\Gamma\rightarrow\mathcal{Y}_{m,0}^{\mathsf{trop}}$ makes a non-zero contribution to the decomposition formula. Then $\Gamma$ contains a vertex $V_{0}$ with the following properties
  1. (1)
    $V_{0}$ is mapped to the origin.
  2. (2)
    $V_{0}$ has $s$ continuable edges
Proof.
There is a distinguished vertex $V_{0}\in V(\Gamma)$ containing the non-tangency marking. We can assume that $h(V_{0})=v_{0}$, for example by imposing the point constraint in the interior of the toric component. By [35, Remark 6.1.1] after potentially subdividing the target further we can assume that $h$ maps vertices to vertices and edges to edges. This allows us to make the gluing argument, as in Proposition 3.3.1, for any $h$. Consequently, there is a virtual dimension constraint on each vertex, which enforces restrictions on $h$ as well as which term in the decomposition of the diagonal contributes to the insertions.
Next we analyse the virtual dimension constraint for $V_{0}$. The virtual dimension of the associated moduli space is $g_{0}+k+k^{\prime}$, where $k+k^{\prime}$ is the valency of $V_{0}$, where $k$ is the number of bounded edges adjacent to $V_{0}$ for which the contribution of the diagonal term consists of $\mathsf{pt}_{E}$ for $V_{0}$. We have that the dimension of the insertions are $g_{0}+s+k$ so that $k^{\prime}=s$.
We will argue all $k^{\prime}$ such edges are continuable. Indeed, if an edge $\gamma_{-1}$ is not continuable then the graph must fork at the vertex of $\gamma_{-1}$ which is not the origin. By Lemma 3.2.5 each branch of the forking can be continued until we find an unbounded edge. Since $\Gamma$ is genus zero each of the unbounded edges found in this way is unique, and it follows that $\Gamma$ has more than $s$ unbounded edges. ∎
We can also exclude the rigid types where a vertex $V\neq V_{0}$ attached to a continuable bounded edge has higher genus by [10, Lemma 14].
Proof of Proposition 3.2.4.
By Lemma 3.2.6 we are left to show that the $k$ bounded edges adjacent to $V_{0}$ for which the diagonal contributes a $\mathsf{pt}_{E}$ to the insertions for $V_{0}$ each attached to a vertex $V_{+}$ which is one valent with image $v_{j}$ for $j\in\{1,\dots,n\}$. If $V_{+}$ is such a vertex then the diagonal must contribute $1$ to the invariant for $V_{+}$ and $\mathsf{pt}_{E}$ for the invariant for $V_{0}$ because the virtual dimension of the moduli space associated to $V_{+}$ is the $g(V_{+})$. Conversely, suppose $V_{+}$ is not a one-valent vertex. By the argument of [12, Lemma 5.8, penultimate paragraph], there is no additional edge connecting $V_{+}$ to a vertex with image $v_{0}$, since there are no additional insertions to make the contribution to that vertex non-zero. Finally, if there is an edge connecting $V_{+}$ to another vertex whose image is not $V_{j}$ then, by Lemma 3.2.5 there would be a sequence of edges which either ends in a cycle, which would give zero contribution due to the $\lambda_{g}$ insertion, or else end in an unbounded edge. But we have already accounted for $s+1$ unbounded edges of $\Gamma$. ∎
Let $k_{\ell ja}$ denote the number of vertices of $\Gamma_{k,\underline{g}}^{s}$ having genus $a$ amongst the $k_{\ell j}$ vertices with curve class $\ell[C_{j}]$ so that
$\sum_{a\geq 0}k_{\ell ja}=k_{\ell j}$
Corollary 3.2.7.
$\mathsf{N}_{g,\vartheta}^{\beta_{p}}=\sum_{k\vdash p}\sum_{\underline{g}}\frac{n_{h^{s}_{k,\underline{g}}}}{|\mathsf{Aut}(h^{s}_{k,\underline{g}})|}\mathsf{N}_{h^{s}_{k,\underline{g}}}=\sum_{k\vdash p}\sum_{\underline{g}}\mathsf{lcm}(|w_{i}(k)|,|\vartheta_{i_{j}}|)\left(\prod_{j=1}^{n}\prod_{\ell\geq 1}\prod_{a\geq 0}\frac{1}{k_{\ell ja}!}\right)\mathsf{N}_{h^{s}_{k,\underline{g}}}$
Proof.
By Proposition 3.2.4 all curves that contribute in the decomposition formula arise from the construction of Definition 3.2.2. The fact that $n_{h^{s}_{k,\underline{g}}}=\mathsf{lcm}(|w_{i}(k)|,|\vartheta_{j}|)$ follows from the fact that $n_{h^{s}_{k,\underline{g}}}$ is the smallest integer such that after scaling by $n_{h^{s}_{k,\underline{g}}}$, $h^{s}_{k,\underline{g}}$ gets integral vertices and integral lengths. The automorphism factor can be computed by observing that those one-valent vertices which have the same curve class and genus can be permuted. ∎
It suffices now to compute $\mathsf{N}_{h_{k,\underline{g}}}^{s}$.
3.3. Vertex contributions
Consider $h_{k,\underline{g}}^{s}:\Gamma_{k,\underline{g}}^{s}\rightarrow\mathcal{Y}_{m,0}^{\mathsf{trop}}$. There are three types of vertices of $\Gamma_{k,\underline{g}}^{s}$
  1. (1)
    The vertex $V_{0}$ which maps via $h_{k,\underline{g}}^{s}$ to $v_{0}$, corresponding to $\overline{Y}_{m}$
  2. (2)
    The vertices $V_{j}$ for $j=1,\dots,s(k)$, which map to vertices corresponding to ~$\tilde{\mathbb{P}}_{i}$
  3. (3)
    The vertices $V^{\prime}_{j}$ for $j=1,\dots,b$, which map to vertices corresponding to ~$\tilde{\mathbb{P}}_{i}$.
3.3.1. Toric contribution
Consider the multiset $\Delta$ consisting of all consisting of $v_{V_{0},E}$ for edges adjacent to $V_{0}$. This multiset consists of $\vartheta_{1},\dots,\vartheta_{s}$ together with the elements of $w(k)=(w_{1}(k),\dots,w_{s(k)}(k))$ and one copy of the zero vector. Associated to this multiset we have $\mathsf{M}_{g_{0},\Delta}$ the moduli space of genus $g_{0}$, $s(k)+s+1$-pointed stable logarithmic maps to $\overline{Y}_{m}$. If $w_{j}(k)=-\ell m_{i}$ let
$\operatorname{\mathsf{ev}}_{j}:\mathsf{M}_{g_{0},\Delta}\rightarrow D_{i}$
denote the evaluation morphism at the $j^{\mathsf{th}}$ marking for $j=1,\dots,s(k)$. Let
$\operatorname{\mathsf{ev}}:\mathsf{M}_{g_{0},\Delta,\Delta_{k}}\rightarrow\overline{Y}_{m}$
denote the evaluation morphism at the last marked point. Define the toric contribution to be
$\mathsf{N}_{g_{0},\Delta,\Delta_{k}}=\int_{[\mathsf{M}_{g_{0},\Delta}]^{\text{\rm vir}}}(-1)^{g_{0}}\lambda_{g_{0}}\prod_{j=1}^{s(k)}\operatorname{\mathsf{ev}}_{j}^{*}(\mathsf{pt}_{D_{i}})\operatorname{\mathsf{ev}}^{*}(\mathsf{pt})\psi^{s-2}$
3.3.2. Non-toric contribution
For each vertex $V_{j}$ for $j=1,\dots,s(k)$ there is an adjacent $E_{j}$ with $v_{V_{j},E_{j}}=\ell m_{i}$ for some $i$. Consider $\mathsf{M}_{a,\ell}(\tilde{\mathbb{P}}_{i})$ the moduli space of $1$-pointed, genus $a$ stable logarithmic maps to ~$\tilde{\mathbb{P}}_{i}$ of degree $\ell[C_{i}]$ with contact order $\ell$ along $D_{m_{i}}$. Define the non-toric contribution to be
$\mathsf{N}_{a}^{\ell\tilde{\mathbb{P}}_{i}}=\int_{[\mathsf{M}_{a,\ell}(\tilde{\mathbb{P}}_{i})]^{\text{\rm vir}}}(-1)^{a}\lambda_{a}$
We will now express $\mathsf{N}_{h_{k,\underline{g}}^{s}}$ as a product over the toric and non-toric contributions. A priori, there are also vertices of type (3), but their will cancel with an overall factor. In  [12, Section 5.5] the author proves a gluing statement at the level of virtual classes on the locus where the curve does not map to the codimension two strata of the pieces of the central fibre. This is sufficient for  [12, Proposition 5.6] as the contribution from the complement of this locus will be killed by the presence of the $\lambda_{g}$ insertion. Similarly to  [26, Section 3.4] this is no longer sufficient for the purposes here. They key point being that  [12, Lemma 5.7] is no longer true in our situation. Instead we follow the same strategy of [26, Section 3.4], by using the gluing formula of [35]. The upshot is the following.
Proposition 3.3.1.
$\mathsf{N}_{h_{k,\underline{g}}^{s}}=\frac{\mathsf{N}_{g_{0},\Delta,\Delta_{k}}}{\mathsf{lcm}(|w_{i}(k)|,|\vartheta_{i_{j}}|)}\left(\prod_{j=1}^{n}\prod_{\ell\geq 1}\ell^{k_{\ell j}}\prod_{a\geq 0}(\mathsf{N}^{\ell\tilde{\mathbb{P}}_{i}}_{a})^{k_{\ell ja}}\right)$
Proof.
We can induct on the bounded edges of $\mathcal{Y}^{\mathsf{trop}}_{m,0}$ and so we can assume that there is only edge. For simplicity, also assume $b=0$.
We wish to turn an integral over $[\mathsf{M}_{h}]^{\text{\rm vir}}$ into one over
$\kappa_{h}^{!}\left([\mathsf{M}_{g_{0},\Delta}]^{\text{\rm vir}}\times\prod_{j=1}^{n}\prod\limits_{\begin{subarray}{c}a\geq 0\\ \ell\geq 1\end{subarray}}([\mathsf{M}_{a,\ell}(\tilde{\mathbb{P}})]^{\text{\rm vir}})^{k_{\ell ja}}\right)$
in the fibre product below, where $\kappa_{h}$ is the diagonal
[…]
If one of the arrows in the above diagram is combinatorially flat, this will follow from  [32, Section 12.3], as the fine and saturated fibre product coincides with the ordinary fibre product. Here we can rewrite the diagram
[…]
so that the right hand vertical arrow is combinatorially flat, as it is a product of $s(k)$ arrows, each combinatorially flat. Consequently, we have that $\mathsf{N}_{h_{k,\underline{g}}^{s}}$ is equal to
$\frac{\prod_{j=1}^{n}\prod_{\ell\geq 1}\ell^{k_{\ell j}}}{\mathsf{lcm}(|w_{i}(k)|)}\int_{[\mathsf{M}_{g_{0},\Delta}]^{\text{\rm vir}}\times\prod_{j,\ell,a}([\mathsf{M}_{a,\ell}(\tilde{\mathbb{P}})]^{\text{\rm vir}})^{k_{\ell ja}}}\left((-1)^{g_{0}}\lambda_{g_{0}}\times\prod_{j,\ell,a}(-1)^{a}\lambda_{a}\times\dots\times(-1)^{a}\lambda_{a}\right)\operatorname{\mathsf{ev}}_{D}^{*}(\kappa_{h})\operatorname{\mathsf{ev}}^{*}({\mathsf{pt}})\psi^{s-2}$
We can decompose the diagonal $\kappa_{h}=\prod_{i=1}^{s(k)}{\mathsf{pt}}_{D}\otimes 1+1\otimes{\mathsf{pt}}_{D}$. Due to the virtual dimension of the moduli spaces, the only summand which contributes is ${\mathsf{pt}}_{D}\otimes 1$ and so we have
$\mathsf{N}_{h_{k,\underline{g}}^{s}}=\frac{\prod_{j=1}^{n}\prod_{\ell\geq 1}\ell^{k_{\ell j}}}{\mathsf{lcm}(|w_{i}(k)|)}\left(\int_{[\mathsf{M}_{g_{0},\Delta}]^{\text{\rm vir}}}(-1)^{g_{0}}\lambda_{g_{0}}\prod_{j=1}^{s(k)}\operatorname{\mathsf{ev}}_{j}^{*}(\mathsf{pt}_{D})\operatorname{\mathsf{ev}}^{*}(\mathsf{pt})\psi^{s-2}\right)\left(\prod_{a\geq 0}\int_{[\mathsf{M}_{a,\ell}(\tilde{\mathbb{P}}_{i})]^{\text{\rm vir}}}(-1)^{a}\lambda_{a}\right)$
which gives the desired formula. In the case where $b\neq 0$, the argument is the same but the contribution from each vertex $V_{i}$ will cancel with the weight of the edge, see  [10, Lemma 15]. ∎
Combining the above with Proposition 3.2.7 gives us the following corollary.
Corollary 3.3.2.
$\mathsf{N}_{g,\vartheta}^{\beta_{p}}=\sum_{k\vdash p}\sum_{\underline{g}}\mathsf{N}_{g_{0},\Delta,\Delta_{k}}\left(\prod_{j\geq 1}\prod_{\ell\geq 1}\ell^{k_{\ell j}}\prod_{a\geq 0}\frac{1}{k_{\ell ja}!}(\mathsf{N}_{a}^{\ell\tilde{\mathbb{P}}_{j}})^{k_{\ell ja}}\right)$
Proposition 3.3.3.
(6) $\sum_{g\geq 0}\mathsf{N}_{g,\vartheta}^{\beta_{p}}u^{2g}=\sum_{k\,\vdash\,p}\left(\sum_{g\geq 0}\mathsf{N}_{g,\Delta,\Delta_{k}}u^{2g+s(k)}\right)\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{\ell^{k_{\ell j}}}{k_{\ell j}!}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{1}{2\sin(\frac{\ell u}{2})}\right)^{k_{\ell j}}$
Proof.
Let $F^{\ell\tilde{\mathbb{P}}_{j}}(u)=\sum_{a\geq 0}\mathsf{N}^{\ell\tilde{\mathbb{P}}_{j}}_{a}u^{2a-1}$ so that
$\left(F^{\ell\tilde{\mathbb{P}}_{j}}(u)\right)^{k_{\ell j}}=\sum_{k_{\ell j}=\sum_{a\geq 0}k_{\ell ja}}\frac{k_{\ell j}!}{\prod_{a\geq 0}k_{\ell ja}}\left(\prod_{a\geq 0}(\mathsf{N}_{a}^{\ell\tilde{\mathbb{P}}_{j}})^{k_{\ell ja}}\right)u^{\sum_{a\geq 0}(2a-1)k_{\ell ja}}.$
Next use that
$2g=2g_{0}+s(k)+(2a-1)\sum_{j=1}^{n}\sum_{\ell\geq 0}\sum_{a\geq 0}k_{\ell ja}$
to conclude
$\sum_{g\geq 0}\mathsf{N}_{g,\vartheta}^{\beta_{p}}u^{2g}=\sum_{k\,\vdash\,p}\left(\sum_{g\geq 0}\mathsf{N}_{g,\Delta,\Delta_{k}}u^{2g+s(k)}\right)\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{\ell^{k_{\ell j}}}{k_{\ell j}!}\left(F^{\ell\tilde{\mathbb{P}}_{j}}(u)\right)^{k_{\ell j}}$
The result now follows from the fact  [9, Lemma 2.20] that
$F^{\ell\tilde{\mathbb{P}}_{j}}(u)=\frac{(-1)^{\ell-1}}{\ell}\frac{1}{2\sin(\frac{\ell u}{2})}.$
Proof of Theorem A.
We now combine Corollary 2.6.2 with Proposition 3.3.3 in order to prove Theorem A. Corollary 2.6.2 tells us that
$[\langle\hat{\vartheta}_{r_{1}},....,\hat{\vartheta}_{r_{s}}\rangle]^{\mathsf{sym}}=\sum_{p}\sum_{k\vdash p}\mathsf{N}^{\Delta,s}_{\mathsf{trop,\Delta_{k}}}(q)\left(\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{1}{k_{\ell j}!}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell j}}\right)\left(\prod_{j=1}^{n}t_{j}^{p_{j}}\right)$
Applying Theorem 1.6.1 to $\mathsf{N}^{\Delta,s}_{\mathsf{trop,\Delta_{k}}}(q)$ tells us that, after identifying $q=e^{iu}$, this is equal to
$\sum_{p}\sum_{k\vdash p}\prod_{j=1}^{n}\prod_{\ell\geq 1}\ell^{k_{\ell j}}\frac{\left(\sum_{g\geq 0}\mathsf{N}_{g,\Delta,\Delta_{k}}u^{2g+s(k)}\right)}{\left((-i)(q^{\frac{1}{2}}-q^{-\frac{1}{2}})\right)^{s(k)}}\left(\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{1}{k_{\ell j}!}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{q^{\frac{1}{2}}-q^{-\frac{1}{2}}}{q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}}\right)^{k_{\ell j}}\right)\left(\prod_{j=1}^{n}t_{j}^{p_{j}}\right).$
Since $s(k)=\sum_{j=1}^{n}\sum_{\ell\geq 1}k_{\ell j}$ and under the change of variables we have
$q^{\frac{\ell}{2}}-q^{-\frac{\ell}{2}}=2i\sin(\frac{\ell u}{2})$
we have that
$[\langle\hat{\vartheta}_{r_{1}},....,\hat{\vartheta}_{r_{s}}\rangle]^{\mathsf{sym}}=\sum_{p}\sum_{k\,\vdash\,p}\left(\sum_{g\geq 0}\mathsf{N}_{g,\Delta,\Delta_{k}}u^{2g+s(k)}\right)\prod_{j=1}^{n}\prod_{\ell\geq 1}\frac{\ell^{k_{\ell j}}}{k_{\ell j}!}\left(\frac{(-1)^{\ell-1}}{\ell}\frac{1}{2\sin(\frac{\ell u}{2})}\right)^{k_{\ell j}}\left(\prod_{j=1}^{n}t_{j}^{p_{j}}\right).$
and so the Theorem follows from Proposition 3.3.3.